# INTERVENOR[Uncaptioned image]: Prompt the Coding Ability of Large Language Models with the Interactive Chain of Repairing

### 19 Jan 2024 v3

@@ 5 Evaluation Results
CodeT5+ Wang et al. (2023) 30.9 - - - - - - - -
InstructCodeT5+ Wang et al. (2023) 35.0 - - - - - - - -
@@
Self-Debug Chen et al. (2023) 73.8 - - - - - - - -
@@
Table 2: Overall Performance of Different Models. We evaluate model effectiveness on code generation and code translation tasks using the $Pass@1$ evaluation metric. The baseline results are borrowed from corresponding papers. More evaluation results on the code translation task are shown in Appendix A.1.
@@ 5.1 Overall Performance
Overall, INTERVENOR outperforms all baselines in all tasks by achieving more than 1% improvements, showing its effectiveness. Compared to our main baseline model GPT-3.5, INTERVENOR achieved about 13% and 4.5% improvements in code generation and code translation tasks, respectively. It illustrates that INTERVENOR has the ability to prompt the coding ability of LLMs by mimicking the human code repairing behavior–iteratively judging, rethinking, and repairing. Notably, INTERVENOR also surpasses Self-Debug and Self-Collaboration models, demonstrating its ability to successfully intervene in the code generation/translation process and guide LLMs to better repair the codes using our chain of repairing mechanism. All these experimental results highlight the generalization ability of INTERVENOR in improving LLMs’ coding ability in different languages.
@@
Table 3: Ablation Studies. We compare different prompting techniques to evaluate the effectiveness of our chain-of-repairing mechanism. All models are built based on GPT-3.5 and evaluated with $Pass@1$.
@@
Table 4: Model Performance of Different Code Learner and Code Teacher Model Pairings. We only conduct one turn code repair in experiments. “N/A” represents the initial output of Code Learner without any intervention from the Code Teacher for code repairing, it reflects the initial generation results of Code Learner.
@@ 5.2 Ablation Studies
Firstly, we evaluate the code generation ability of LLMs using different code generation/repairing prompting methods. As shown in Table 3, we compare CoT Kojima et al. (2023), Few-Shot Chen et al. (2021), and Few-Shot CoT Wei et al. (2023) models in experiments, which prompt LLMs to better generate codes. These models try to generate natural language as the chain of coding thought (CoT) or provide some instances to demonstrate the coding task (Few-Shot). Then we compare different methods to generate the code repair instruction, including Self-Refine Madaan et al. (2023), Error Msgs, and Chain-of-Repairing (CoR). Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs and CoR incorporate the code error messages from compilers. Msgs directly uses error messages to guide the code repair process.
@@
The experimental results show that code repairing is more effective than directly prompting LLMs to generate codes. Even though we conduct different methods to directly prompt LLMs to generate better codes, we achieve only 3.5% improvements. On the contrary, the code repairing methods improve the quality of generated codes by achieving more than 9.7% improvements with only single-turn repairing. Then we compare different code repairing methods. Both Error Msgs and CoR thrive on the feedback from code compilers and achieve more than 1.3% improvements than Self-Refine, demonstrating that compilers can provide valuable signals to help LLMs better recognize the code bugs. Notably, CoR achieves the best performance among all code repairing models, which illustrates its effectiveness in teaching LLMs for code repairing. The CoR mechanism uses the error messages from code compilers to prompt LLMs, aiming to rethink the reasons for making errors and generate the instructions for repairing.
@@
Then, as shown in Table 4, we investigate the impact of using different LLMs to build Code Learner and Code Teacher. When we use the same LLM to build the Code Learner, the performance of INTERVENOR improves as the capabilities of the Code Teacher are enhanced, showing the effectiveness of our CoR mechanism. This indicates that stronger models can conduct a more in-depth analysis of erroneous code and provide more accurate and relevant suggestions for code repair. Similarly, by keeping the same LLM to build the Code Teacher, a stronger Code Learner also conducts better code repair results. This indicates that a more powerful Code Learner has a stronger instruct following ability to better comprehend the CoR provided by the Code Teacher for code repair.
@@
Figure 3: The Impact of Different Code Repairing Turns. HumanEval, MBPP, and HumanEval-X (HEX) are used to evaluate our INTERVENOR model.
@@
Figure 4: Code Repairing Performance on the CodeError Dataset. We repair the error codes with one single turn. The codes are divided into two groups to evaluate the code repairing effectiveness, including Assertion Errors and Others (AttributeError, NameError, RecursionError, SyntacError and TypeError).
5.3 Effectiveness of INTERVENOR in Different Testing Scenarios
In this subsection, we delve deeper into exploring the effectiveness of INTERVENOR in two testing scenarios: 1) validating the impact of different code repairing turns, and 2) evaluating the code repairing effectiveness on different code error types.
Figure 5: Case Studies. We provide two cases that showcase the effectiveness of the Chain of Repairing generated by INTERVENOR when fixing AttributeError and AssertionError, respectively.
As shown in Figure 3, the code generation performance is significantly improved during the iteratively repairing. After three turns, the INTERVENOR achieves almost the best performance on HumanEval and HumanEval-X datasets, showing the efficiency of our interactive chain-of-repairing mechanism. For a more difficult dataset MBPP, INTERVENOR achieves 30% improvements and still has some room to achieve further improvements, demonstrating its advantages in dealing with more difficult and realistic coding tasks.
Then we show the code repairing effectiveness on different error types in Figure 4. We use different prompt methods to stimulate LLMs to repair code errors and show their effectiveness on different types of code errors. Overall, INTERVENOR doubles the number of corrected code examples of baseline models on both AssertionError and other error types, showing its effectiveness in repairing the code errors. On the one hand, INTERVENOR significantly outperforms other methods on the AssertionError task. It illustrates that INTERVENOR can provide more precise guidance and identify errors with the help of the bug report derived from the failed testing case. On the other hand, INTERVENOR also shows a strong ability to correct other code errors. Our chain-of-repairing method also thrives on the error messages of the compiler, breaks the cognitive inertia of LLMs, and identifies the specific code error for repairing. All these phenomena show that the quality of code execution feedback is critical in repairing codes.
@@ 5.4 Case Studies
Finally, we select two case studies in Figure 5 to demonstrate the effectiveness of INTERVENOR. We also compare INTERVENOR with Self-Refine Madaan et al. (2023), which prompts LLMs to execute codes, recognize bugs, and generate instructions by themselves.
Overall, the feedback from compilers indeed helps to improve the accuracy of code repairing by providing more valuable instructions. In the first case, Self-Refine fails to fix the AttributeError and adds the “all()” function in codes. On the contrary, INTERVENOR successfully fixes the code, showing its effectiveness. It accurately analyzes the reason for bugs “call the islower() method on an integer object, which is not possible” and also provides a solution by suggesting to “check if the key is a string before calling the islower() method”. In the second case, the instruction provided by Self-Refine thinks the code seems to be correct and does not offer definite solutions, showing the cognitive inertia in debug–It is hard to debug the code written by ourselves. INTERVENOR shows its effectiveness in directly generating the reason of bug “the function is not handling the case where the input list has only one element correctly”. More cases are shown in Appendix A.3.
@@ 6 Conclusion
This paper proposes INTERVENOR, which regards the feedback of compilers as the INTERVENOR to facilitate the agent collaboration of rethinking and repairing. Our experiments show that INTERVENOR outperforms previous code generation/repairing models by accurately diagnosing the buggy codes with the CoR mechanism.
@@
The performance of INTERVENOR largely depends on the capabilities of the foundation model, requiring the model to have the ability to analyze code errors and repair code, which poses higher demands on the model’s capabilities. The gains brought by INTERVENOR with models of inferior capabilities might be limited. Moreover, for some codes with only logical errors, the compiler requires the assistance of test cases to provide feedback. The quality and coverage of test cases are crucial in this scenario. Inadequate or poorly designed test cases may not be sufficient to detect or help rectify these logical errors effectively. Therefore, the effectiveness of INTERVENOR also hinges on the robustness and comprehensiveness of the test cases used in conjunction with it.
@@ Appendix A Appendix
A.1 Code Translation Results.
@@
A.2 More Details of the CodeError Benchmark
In our experiments, we build the CodeError benchmark to evaluate the code repair capabilities of LLMs, which facilitates the research on the code repairing task. The CodeError dataset is collected from basic programming problems, data analysis problems, and programming competition problems. In this section, we show more detailed data information about the CodeError benchmark.
Figure 6: An Example of the CodeError Dataset. The CodeError benchmark asks LLMs to fix buggy codes and then evaluate whether the fixed code meets the requirements specified in the task description and passes all test cases. During the process of code repairing, LLMs can utilize error messages to repair codes.
@@ A.2 More Details of the CodeError Benchmark
Table 5: Code Translation Performance. We evaluate the code translation effectiveness among different program languages, including Python, C++, Java, and JavaScript (JS). We report the results of INTERVENOR, which only repair code with a single turn. All evaluation results are evaluated with $Pass@1$.
@@
Table 6: A Comparison between CodeError and Other Code Repair Benchmarks. Size only represents the size of the test set. Avg.TC indicates the average number of test cases per problem. CE indicates compilation errors (e.g., SyntaxError). Error Msg indicates whether the buggy code contains detailed error information, such as the line of incorrect code, the reasons for the error, etc. Category represents the scope covered by the buggy code, CodeError covers basic programming problems, data analysis problems, and programming competition problems.
@@
CodeError Examples. As shown in Figure 6, each example in CodeError consists of a programming task description, a buggy code snippet, error messages, and test cases to evaluate the code correctness of the generated codes. If the fixed/generated code meets the requirements specified in the task description and passes all test cases, it proves that the repaired code is correct. During the process of code repairing, LLMs need to fix the buggy code and can also use the error messages from the buggy code for assistance. We provide more examples that are sampled from the basic programming problems, programming competition problems, and data analysis problems in Figure 7, Figure 8, and Figure 9, respectively.
Data Collection. CodeError is collected from different coding problems, making the dataset diverse and reliable.
To ensure the diversity of CodeError, we collect basic programming problems, data analysis problems, and programming competition problems from HumanEval, MBPP, DS-1000, APPS, and the programming contest site AtCoder. We conduct code generation experiments using GPT-3.5 and CodeLlama on the first four datasets, preserving the generated codes that contain errors. For the programming contest site AtCoder, we crawl real-world user-submitted buggy codes to ensure the reliability of CodeError. The diverse data sources allow us to build a comprehensive and robust dataset to estimate the code repairing ability of LLMs.
Data Statistics. The data statistics of CodeError are shown in Table  7.
@@
Similar to FixEval Haque et al. (2023), we utilize test cases to assess the functional correctness of the repaired code. Additionally, in comparison to other benchmarks Yasunaga and Liang (2021); Huq et al. (2022); Lu et al. (2021); Haque et al. (2023), we have two notable features. 1) We provide detailed error information for erroneous code, including error code localization, the cause of the error, error type, and more. 2) CodeError covers a variety of problem types, including basic programming problems, data analysis problems, and programming competition problems, enabling a more comprehensive benchmark to evaluate the code repairing capabilities of LLMs.
A.3 Additional Case Studies
In this subsection, we sample some cases from CodeError to demonstrate the effectiveness of the Chain of repairing (CoR) mechanism.
As shown in Figure 10, in the first case, there is a simple NameError in the buggy code, which indicates that ‘hashlib’ is not defined. We can see that the INTERVENOR recognizes hashlib as a Python package and provides a solution: “import the ‘hashlib’ module at the beginning of the code”. In the second case, the error in the code is more subtle, involving an operation between ‘int’ and ‘list’. The CoR explicitly states that this is a “valid operation” and provides a solution: “check if the current element in the list is a list or not before adding it to the sum”. With the help of the accurate instruction, INTERVENOR successfully repairs the code. In Figure 11, we demonstrate that INTERVENOR can effectively address the boundary issues, which are often quite tricky. INTERVENOR accurately identifies the code error in the loop condition and suggests a modification method: “modify the loop to iterate up to $n-1$ instead of $n$”. In Figure 12, we demonstrate that INTERVENOR can fix multiple error types in the code by iteratively repairing. Over two iterations, INTERVENOR identifies and accurately modifies the ValueError and TypeError in the code.
All these cases demonstrate the effectiveness of the chain of repairing (CoR) mechanism and highlight the effectiveness of CoR in identifying errors and providing appropriate solutions for code repairing.

### 20 Feb 2024 v4

@@ 5 Evaluation Results
CodeT5+ Wang et al. (2023b) 30.9 - - - - - - - -
InstructCodeT5+ Wang et al. (2023b) 35.0 - - - - - - - -
@@
Self-Debug Chen et al. (2023b) 73.8 - - - - - - - -
@@
Table 2: Overall Performance of Different Models. We evaluate model effectiveness on code generation and code translation (HumanEval-X dataset) tasks using the Pass@$1$ evaluation metric. The baseline results are borrowed from corresponding papers. More evaluation results on the code translation task are shown in Appendix A.2.
@@ 5.1 Overall Performance
Overall, INTERVENOR outperforms all baselines in all tasks by achieving more than 1% improvements, showing its effectiveness. Compared to our main baseline model GPT-3.5, INTERVENOR achieved about 18% and 4.3% improvements in code generation and code translation tasks, respectively. It illustrates that INTERVENOR has the ability to prompt the coding ability of LLMs by mimicking the human code repair behavior–iteratively judging, rethinking, and repairing. Notably, INTERVENOR also surpasses Self-Debug and Self-Collaboration models, demonstrating its ability to successfully intervene in the code generation/translation process and guide LLMs to better repair the codes using our chain of repair mechanism. All these experimental results highlight the generalization ability of INTERVENOR in improving LLMs’ coding ability in different languages.
@@
Table 3: Ablation Studies. We compare different prompting techniques to evaluate the effectiveness of our Chain-of-Repair (CoR) mechanism. All models are built based on GPT-3.5 and evaluated with Pass@$1$.
@@
Table 4: Model Performance of Different Code Learner and Code Teacher Model Pairings. We only conduct one turn code repair in experiments. “N/A” represents the initial output of Code Learner without any intervention from the Code Teacher for code repair, it reflects the initial generation results of Code Learner.
@@ 5.2 Ablation Studies
Firstly, we evaluate the code generation ability of LLMs using different code generation/repair prompting methods. As shown in Table 3, we compare CoT Kojima et al. (2022), Few-Shot Chen et al. (2021), and Few-Shot CoT Wei et al. (2022) models in experiments, which prompt LLMs to better generate codes. These models try to generate natural language as the chain of coding thought (CoT) or provide some instances to demonstrate the coding task (Few-Shot). Then we compare different methods to generate the code repair instruction, including Self-Refine Madaan et al. (2023), Error Msgs, and Chain-of-Repair (CoR). Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs and CoR incorporate the code error messages from compilers. Msgs directly uses error messages to guide the code repair process.
@@
The experimental results show that code repair is more effective than directly prompting LLMs to generate codes. Even though we conduct different Few-shot and CoT methods to directly prompt LLMs, we only achieve 3.5% improvements, which shows that it is difficult for LLMs to generate correct codes without repairing. On the contrary, the code repair methods improve the quality of generated codes by achieving more than 9.7% improvements with only single-turn repair. Both Error Msgs and CoR thrive on the feedback from code compilers and achieve more than 1.2% improvements than Self-Refine, demonstrating that compilers can provide valuable signals to help LLMs better recognize the code bugs. Notably, CoR achieves the best performance among all code repair models, illustrating its effectiveness in guiding LLMs for code repair. CoR uses the error messages from code compilers to prompt LLMs, aiming to rethink the reasons for making errors and generate the instructions for repairing.
@@
Then, as shown in Table 4, we investigate the impact of using different LLMs to build Code Learner and Code Teacher. When we use the identical LLM to build the Code Learner, the performance of INTERVENOR is improved with the enhanced capabilities of the Code Teacher, underscoring the efficacy of our CoR mechanism. This indicates that stronger models can conduct a more in-depth analysis of erroneous code and provide more accurate suggestions for code repair. Similarly, by keeping the same LLM to build the Code Teacher, a stronger Code Learner also conducts better code repair results. This indicates that a more powerful Code Learner has a stronger instruction-following ability to better comprehend the CoR provided by the Code Teacher.
@@
Figure 3: The Impact of Different Code Repair Turns. HumanEval, MBPP, and HumanEval-X (HEX) are used to evaluate our INTERVENOR model.
5.3 Effectiveness of INTERVENOR in Different Testing Scenarios
In this subsection, we delve deeper into exploring the effectiveness of INTERVENOR in two testing scenarios: 1) validating the impact of different code repair turns, and 2) evaluating the code repair effectiveness on different code error types.
As shown in Figure 3, the code generation performance is significantly improved during the iteratively repairing. After three turns, the INTERVENOR achieves almost the best performance on HumanEval and HumanEval-X datasets, showing the efficiency of our interactive chain-of-repair mechanism. For a more difficult dataset MBPP, INTERVENOR achieves 30% improvements and still has some room to achieve further improvements, demonstrating its advantages in dealing with more difficult and realistic coding tasks.
@@ 5.3 Effectiveness of INTERVENOR in Different Testing Scenarios
Figure 4: Code Repair Performance on the CodeError Dataset. We repair the error codes with one single turn. The codes are divided into two groups to evaluate the code repair effectiveness, including Assertion Errors and Others (AttributeError, NameError, RecursionError, SyntacError and TypeError).
Figure 5: Case Studies. We provide two cases that showcase the effectiveness of the Chain-of-Repair (CoR) generated by INTERVENOR when fixing AttributeError and AssertionError, respectively.
Then we show the code repair effectiveness on different error types in Figure 4. We use different prompt methods to stimulate LLMs to repair code errors and show their effectiveness on different types of code errors. Overall, INTERVENOR doubles the number of corrected code examples of baseline models, showing its effectiveness in repairing code errors. On the one hand, INTERVENOR significantly outperforms other methods on the AssertionError repairing task. It illustrates that INTERVENOR can provide more precise guidance and identify errors with the help of the bug report derived from the failed testing case. On the other hand, INTERVENOR also shows a strong ability to correct other code errors. Our Chain-of-Repair (CoR) method also thrives on the error messages of the compiler, breaks the cognitive inertia of LLMs, and identifies the specific code error for repairing. All these phenomena show that the quality of code execution feedback is critical in repairing codes.
@@ 5.4 Case Studies
Finally, we show two cases in Figure 5 to demonstrate the effectiveness of INTERVENOR. We compare INTERVENOR with Self-Refine Madaan et al. (2023), which prompts LLMs themselves to execute codes, recognize bugs, and repair codes.
Overall, the feedback from compilers indeed helps to improve the accuracy of repaired code by providing more valuable instructions. In the first case, Self-Refine fails to fix the AttributeError and adds the “all()” function in codes. On the contrary, INTERVENOR successfully fixes the code, showing its effectiveness. It accurately analyzes the reason of bugs “call the islower() method on an integer object, which is not possible” and also provides a solution by suggesting to “check if the key is a string before calling the islower() method”. In the second case, the instruction provided by Self-Refine thinks the code seems to be correct and does not offer definite solutions, showing the cognitive inertia in debug–It is hard to debug the code written by ourselves. INTERVENOR shows its effectiveness in directly generating the reason for the bug: “the function is not handling the case where the input list has only one element correctly”. More cases are shown in Appendix A.4.
@@ 6 Conclusion
This paper proposes INTERVENOR, which regards the feedback of compilers as the INTERVENOR to facilitate the agent collaboration of code error diagnose and repair. Our experiments show that INTERVENOR outperforms previous code generation/repair models by generating accurate code repair solutions and instructions.
@@
The effectiveness of INTERVENOR heavily depends on the capabilities of the foundation model, necessitating its proficiency in code error analysis and repair. The efficacy of INTERVENOR may be constrained when using less proficient LLMs to build the agents. Furthermore, in cases where codes exhibit solely logical errors, the compiler relies on test cases to furnish feedback. The quality and coverage of test cases are crucial in this scenario. Inadequate or poorly designed test cases may not be sufficient to detect or help rectify these logical errors effectively. Hence, the efficacy of INTERVENOR is intricately linked to the robustness and comprehensiveness of the test cases.
Ethics Statement
In conducting this research and presenting the findings in this paper, we have upheld rigorous ethical standards throughout the entirety of the process. Our experiments are conducted on commonly used datasets from previous works. Additionally, for our newly constructed dataset, we have also reached out to relevant platforms, and they have granted us permission to use the data for scientific research purposes. Besides, the code submitted to the Atcoder website has undergone stringent ethical review, with very few instances where ethical concerns arise. Furthermore, despite the risk of large language models (LLMs) generating toxic data, by providing clear role-playing instructions and only allowing LLMs to complete code-related tasks, we believe the model’s output will not produce harmful content.
@@ Appendix A Appendix
A.1 License
For all datasets in our experiments, HumanEval uses the MIT License, MBPP uses the CC-BY-4.0 License, and HumanEval-X uses the Apache License 2.0. All of these licenses allow their data for academic use.
A.2 Code Translation Results
@@
A.3 More Details of the CodeError Benchmark
In our experiments, we build the CodeError benchmark to evaluate the code repair capabilities of LLMs, which facilitates the research on the code repair task. The CodeError dataset is collected from basic programming problems, data analysis problems, and programming competition problems. In this section, we show more detailed data information about the CodeError benchmark.
Figure 6: An Example of the CodeError Dataset. The CodeError benchmark asks LLMs to fix buggy codes and then evaluate whether the fixed code meets the requirements specified in the task description and passes all test cases. During the process of code repair, LLMs can utilize error messages to repair codes.
@@ A.3 More Details of the CodeError Benchmark
Table 5: Code Translation Performance on Humaneval-X. We evaluate the code translation effectiveness among different program languages, including Python, C++, Java, and JavaScript (JS). We report the results of INTERVENOR, which only repairs codes with a single turn. All evaluation results are evaluated with Pass@$1$.
@@
Table 6: A Comparison between CodeError and Other Code Repair Benchmarks. Size only represents the size of the test set. Avg.TC indicates the average number of test cases per problem. CE indicates compilation errors (e.g., SyntaxError). Error Msg indicates whether the buggy code contains detailed error information, such as the line of incorrect code, the reasons for the error, etc. Category represents the scope covered by the buggy code, and CodeError covers basic programming problems, data analysis problems, and programming competition problems.
@@
CodeError Examples. As shown in Figure 6, each example in CodeError consists of a programming task description, a buggy code snippet, error messages, and test cases to evaluate the code correctness of the repaired codes. If the fixed/generated code meets the requirements specified in the task description and passes all test cases, it proves that the repaired code is correct. During the process of code repair, LLMs need to fix the buggy code and are also able to use the error messages from the buggy code for assistance. We provide more examples that are sampled from the basic programming problems, programming competition problems, and data analysis problems in Figure 7, Figure 8, and Figure 9, respectively.
Data Collection. CodeError is collected from various coding problems, making the dataset diverse and reliable.
To ensure the diversity of CodeError, we collect basic programming problems, data analysis problems, and programming competition problems from HumanEval, MBPP, DS-1000, APPS, and the programming contest site AtCoder. We conduct code generation experiments using GPT-3.5 on the first four datasets, preserving the generated codes that contain errors. For the programming contest site AtCoder, we crawl real-world user-submitted buggy codes to ensure the reliability of CodeError. The diverse data sources allow us to build a comprehensive and robust dataset to estimate the code repair ability of LLMs.
Data Statistics. The data statistics of CodeError are shown in Table 7.
@@
Similar to FixEval Haque et al. (2023), we utilize test cases to assess the functional correctness of the repaired code. Additionally, in comparison to other benchmarks Yasunaga and Liang (2021); Huq et al. (2022); Lu et al. (2021); Haque et al. (2023), we have two notable features:
  • We provide detailed error information for erroneous codes, including error code localization, the cause of the error, error type, and more.
  • CodeError covers a variety of problem types, including basic programming problems, data analysis problems, and programming competition problems, enabling a more comprehensive benchmark to evaluate the code repair capabilities of LLMs.
A.4 Additional Case Studies
In this subsection, we sample some cases from CodeError to demonstrate the effectiveness of the Chain-of-Repair (CoR) mechanism.
As shown in Figure 10, in the first case, there is a simple NameError in the buggy code, which indicates that ‘hashlib’ is not defined. We can see that the INTERVENOR recognizes hashlib as a Python package and provides a solution: “import the ‘hashlib’ module at the beginning of the code”. In the second case, the error in the code is more subtle, involving an operation that should check the data structures of ‘int’ and ‘list’. The CoR explicitly states that this is a “valid operation” and provides a solution: “check if the current element in the list is a list or not before adding it to the sum”. With the help of accurate instruction, INTERVENOR successfully repairs the code.
Additionally, in Figure 11, we demonstrate that INTERVENOR can effectively address boundary issues, which are often quite tricky. INTERVENOR accurately identifies the code error in the loop condition and suggests a modification method: “modify the loop to iterate up to $n-1$ instead of $n$”. In Figure 12, we demonstrate that INTERVENOR can fix multiple error types in the code by iteratively repairing. Over two iterations, INTERVENOR identifies and accurately modifies the ValueError and TypeError in the code.
All these cases demonstrate the effectiveness of the Chain-of-Repair (CoR) mechanism and highlight the effectiveness of CoR in identifying errors and providing appropriate solutions for code repair.